FInd the slope
y=x-1/2

Answer by rapaljer(4671)

In the form y=mx+b, the slope is always the coefficient of x. In this case,
y=x-1/2, it really means y=1x-1/2, so the coefficient of x is 1. Therefore, slope =1.

Dr. Robert J. Rapalje, Retired